# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 3.3 Computation of the Hessian
The second-order derivative of EOT distance $\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to the source data $\boldsymbol{Y}$ is given by the fourth order tensor $\mathcal{T}\in\mathbb{R}^{M\times d\times M\times d}$
$\mathcal{T}_{ktsl}=\frac{1}{\epsilon}\sum_{i,j=1}^{M+N}\mathcal{R}_{ikt}H^{\dagger}_{ij}\mathcal{R}_{jsl}+\mathcal{E}_{ktsl},\text{ for }k,s=1,\dots,M\text{ and }t,l=1,\dots,d,$ (36)
where $H^{\dagger}$ is the Moore–Penrose inverse of matrix $H(\Pi^{*})\in\mathbb{R}^{(M+N)\times(M+N)}$ defined in (2), and $\mathcal{R}=\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbbm{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]\in\mathbb{R}^{(M+N)\times M\times d}$ is the right-hand-side third-order tensor defined in (34). The fourth-order tensor $\mathcal{E}\in\mathbb{R}^{M\times d\times M\times d}$ is defined as
$\mathcal{E}_{ktsl}=\begin{cases}\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\left(\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}\right)_{tl}-\frac{1}{\epsilon}\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{t}}\cdot\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{l}}\right),&\text{if}k=s\\ 0,&\text{Otherwise}\end{cases}$ (37)
@@
$\frac{d^{2}\mathrm{OT}_{\epsilon}(C,\mu,\nu)}{d\boldsymbol{y}_{s}d\boldsymbol{y}_{k}}=\begin{cases}\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{s}},&\text{if }k\neq s,\\ \sum_{j=1}^{N}\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}\Pi_{kj}^{*}+\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{k}},&\text{if }k=s.\end{cases}$ (38)
@@
$\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{s}}=\sum_{j=1}^{N}\frac{\Pi^{*}_{kj}}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\left(\frac{\partial f^{*}_{k}}{\partial\boldsymbol{y}_{s}}+\frac{\partial g^{*}_{j}}{\partial\boldsymbol{y}_{s}}\right)$
$=\left[\begin{array}[]{ccccccccc}\boldsymbol{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right]^{\top}\frac{H^{\dagger}}{\epsilon}\left[\begin{array}[]{ccccccccc}\boldsymbol{e}_{s}(B_{s}\mathbbm{1}_{N})^{\top}\\ B_{s}^{\top}\end{array}\right],$
where we use the fact that $\left[\begin{array}[]{ccccccccc}\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{s}}\\ \frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{s}}\end{array}\right]$ is the solution of the linear system (26), so it can be expressed as $\left[\begin{array}[]{ccccccccc}\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{s}}\\ \frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{s}}\end{array}\right]=H^{\dagger}\left[\begin{array}[]{ccccccccc}\mathbf{e}_{s}(B_{s}\mathbbm{1}_{N})^{\top}\\ B_{s}^{\top}\end{array}\right]$. For $k=s$, the term is expanded as follows
$\sum_{j=1}^{N}\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}\Pi_{kj}^{*}+\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{k}}$
$\,=\sum_{j=1}^{N}\frac{\Pi^{*}_{kj}}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\left(\frac{\partial f^{*}_{k}}{\partial\boldsymbol{y}_{k}}+\frac{\partial g^{*}_{j}}{\partial\boldsymbol{y}_{k}}\right)+\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}-\frac{1}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)\right)$
$\,=\left[\begin{array}[]{ccccccccc}\boldsymbol{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right]^{\top}\frac{H^{\dagger}}{\epsilon}\left[\begin{array}[]{ccccccccc}\boldsymbol{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right]+\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}-\frac{(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}})^{\top}(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}})}{\epsilon}\right).$
@@
Given the cost matrix $C$ and the optimal coupling matrix $\Pi^{*}$ from Sinkhorn algorithm, the Hession tensor is analytically calculated once the large linear system (34) is numerically solved. Unlike the previous approaches, the solution of the linear system is no longer needed to be differentiated in order to obtain the second-order derivatives, so our direct analytical Hessian expression could significantly speed up the computation with less memory burden. Similar to the case of the first-order derivative, the analytical expression for the Hessian is generic and applicable to any choice of cost function. It does not require $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ to be uniform either. If the Sinkhorn iterations stops early with $\widehat{\Pi}$ being suboptimal, then the Hessian expression (36) is still exact for $\frac{d^{2}\mathrm{OT}_{\epsilon}(C,{\tilde{\boldsymbol{\mu}}},{\tilde{\boldsymbol{\nu}}})}{d\boldsymbol{Y}^{2}}$ with suboptimal coupling matrix $\widehat{\Pi}$ and associated marginals $\tilde{\boldsymbol{\mu}}$ and $\tilde{\boldsymbol{\nu}}$, where $\sum_{j}\widehat{\Pi}_{ij}=\widetilde{\mu}_{i}$ and $\sum_{i}\widehat{\Pi}_{ij}=\widetilde{\nu}_{j}$.
With the analytical expression of Hessian, the following result show that the sum of the first index of the Hessian tensor is only dependent on the marginal probability vector $\boldsymbol{\mu}$. Later, it will be used as a marginal error to verify the accuracy of Hessian in the numerical implementation.
@@
If $C_{kj}=\|\boldsymbol{y}_{k}-\boldsymbol{y}_{j}^{*}\|_{2}^{2}$ for each $k,j$, then $\sum_{k=1}^{M}\mathcal{T}_{k\cdot s\cdot}=2\mu_{s}\mathbb{I}_{d}$.
@@
Since the cost is square distance $C_{kj}=\|\boldsymbol{y}_{k}-\boldsymbol{y}_{j}^{*}\|_{2}^{2}$, then $\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}=2(\boldsymbol{y}_{k}-\boldsymbol{y}_{j}^{*})$ and $\frac{\partial^{2}C_{kj}}{\partial\boldsymbol{y}_{k}^{2}}=2\mathbb{I}_{d}$. The first term in the second-order derivative (38) expands as
$\sum_{k=1}^{M}\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{s}}$ $=2\sum_{k=1}^{M}\boldsymbol{y}_{k}^{\top}\sum_{j=1}^{N}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{s}}-2\sum_{j=1}^{N}(\boldsymbol{y}_{j}^{*})^{\top}\sum_{k=1}^{M}\frac{\partial\Pi_{kj}^{*}}{\partial\boldsymbol{y}_{s}}=\mathbb{0}_{d\times d}.$
The last equal sign is because that $\sum_{j}\Pi^{*}_{kj}=\mu_{k}$ and $\sum_{k}\Pi^{*}_{kj}=\nu_{j}$, as well as $\sum_{j}\left(\frac{\partial\Pi^{*}_{kj}}{\partial\boldsymbol{y}_{s}}\right)=\frac{\partial\sum_{j}\Pi^{*}_{kj}}{\partial\boldsymbol{y}_{s}}$ and $\sum_{k}\left(\frac{\partial\Pi^{*}_{kj}}{\partial\boldsymbol{y}_{s}}\right)=\frac{\partial\sum_{k}\Pi^{*}_{kj}}{\partial\boldsymbol{y}_{s}}$.
The remaining term is $\sum_{j=1}^{N}\frac{\partial^{2}C_{sj}}{\partial\boldsymbol{y}_{s}^{2}}\Pi^{*}_{sj}=2\mathbb{I}_{d}\sum_{j=1}^{N}\Pi^{*}_{sj}=2\mu_{s}\mathbb{I}_{d}.$ Thus we proved the statement. ∎
@@ 3.4 Solve the linear systems with truncated SVD
A major challenge in computing the gradient of Sinkhorn distance in (35) and the Hessians of Entroy-regularized distance in (36) is that pseudo-inverse $H(\Pi^{*})^{\dagger}$ can severely amplify the rounding errors or early stopping errors when the matrix $H(\Pi^{*})$ is ill-conditioned. The pseudo-inverse comes from the solutions to the linear systems (26) and (34) for computation of first- and second- order derivatives. Thus, instead of directly using $H(\Pi^{*})^{\dagger}$, it is important to study the spectrum of the matrix $H(\Pi^{*})$ and regularize properly when the linear systems are ill-posed.
@@
Analytical and empirical results in the next section show that the $H$-matrix is often ill-conditioned when the Sinkhorn regularization parameter $\epsilon$ is small or when the optimal coupling matrix $\Pi^{*}$ is close to a permutation. Such ill-conditioned $H$-matrices result in numerically unstable solutions when $\Pi^{*}$ is slightly perturbed to $\widehat{\Pi}$ due to the early stopping of the Sinkhorn iterations. This is also the exact reason that the previous implicit differentiation approach fails due to numerically instability.
@@
We tackle the potential ill-posedness by truncated Singular value decomposition (TSVD). We truncate the H-matrix’s spectrum up to the $K$-th largest eigenvalue, with $K=\max\{j:\frac{\lambda_{j}}{\lambda_{1}}>\alpha\}$, where $\{\lambda_{j}\}_{j=1}^{M+N}$ are the eigenvalues of $H(\Pi^{*})$ in descending order. In practice, we use the LAPACK’s DGELSD algorithm building in the least-square solver [4] and set $\alpha=10^{-10}$. The algorithm for gradient and Hessian computation is summarized in Algorithm 1.
@@
Algorithm 1 Computation of gradient and Hessian of EOT distance $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to the source data $\boldsymbol{Y}$.
@@
Input: Optimized $\Pi^{*}$, cost matrix $C$, entropic regularization strength $\epsilon$, source data $\boldsymbol{Y}$ and singular value threshold $\alpha$.
@@
Output: Gradient with respect to $\boldsymbol{Y}$: $\frac{d\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}\in\mathbb{R}^{M\times d}$, Hessian with respect to $\boldsymbol{Y}$: $\mathcal{T}\in\mathbb{R}^{M\times d\times M\times d}$
1:Compute the marginal probability vector $\boldsymbol{\mu}\in\mathbb{R}^{M}$ and $\boldsymbol{\nu}\in\mathbb{R}^{N}$: $\boldsymbol{\mu}\leftarrow\Pi^{*}\mathbbm{1}_{N}$ and $\boldsymbol{\nu}\leftarrow(\Pi^{*})^{\top}\mathbbm{1}_{M}$.
@@
2:Compute matrix $H\in\mathbb{R}^{(M+N)\times(M+N)}$: $H\leftarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\boldsymbol{\mu})&\Pi^{*}\\ (\Pi^{*})^{\top}&\text{diag}(\boldsymbol{\nu})\end{array}\right]$.
@@
3:Compute third-order tensor $\mathcal{B}\in\mathbb{R}^{M\times d\times N}$: $\mathcal{B}_{ksj}\leftarrow\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{s}}\Pi^{*}_{kj}$, compute third order tensor $\mathcal{R}\in\mathbb{R}^{(M+N)\times M\times d}$: $\mathcal{R}\leftarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbbm{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]$ and compute the fourth-order tensor $\mathcal{E}$ in (37).
@@
4:Compute the gradient: $\frac{d\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}=\mathcal{B}\cdot\mathbbm{1}_{N}$.
@@
6:Approximate $\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]$: $\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]\leftarrow U_{K}\Lambda_{K}^{-1}U_{K}^{\top}\mathcal{R}$.
7:Compute the Hessian $\mathcal{T}$: $\mathcal{T}_{ktsl}\leftarrow\frac{1}{\epsilon}\sum_{i=1}^{M+N}\mathcal{R}_{ikt}\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]_{isl}+\mathcal{E}_{ktsl}$.

### v2

@@ 3.3 Computation of the Hessian
The second-order derivative of EOT distance $\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ with respect to the source data $\mathbf{Y}$ is given by the fourth order tensor $\mathcal{T}\in\mathbb{R}^{M\times d\times M\times d}$
$\mathcal{T}_{ktsl}=\frac{1}{\epsilon}\sum_{i,j=1}^{M+N}\mathcal{R}_{ikt}(\mathbf{H}^{*})^{\dagger}_{ij}\mathcal{R}_{jsl}+\mathcal{E}_{ktsl},\text{ for }k,s=1,\dots,M\text{ and }t,l=1,\dots,d,$ (34)
where $(\mathbf{H}^{*})^{\dagger}$ is the Moore–Penrose inverse of matrix $\mathbf{H}^{*}\in\mathbb{R}^{(M+N)\times(M+N)}$ defined in (2), and $\mathcal{R}=\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbb{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]\in\mathbb{R}^{(M+N)\times M\times d}$ is the right-hand-side third-order tensor defined in (32). The fourth-order tensor $\mathcal{E}\in\mathbb{R}^{M\times d\times M\times d}$ is defined as
$\mathcal{E}_{ktsl}=\begin{cases}\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\left(\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}\right)_{tl}-\frac{1}{\epsilon}\frac{\partial C_{kj}}{\partial(\mathbf{y}_{k})_{t}}\cdot\frac{\partial C_{kj}}{\partial(\mathbf{y}_{k})_{l}}\right),&\text{if }k=s\\ 0,&\text{Otherwise}\end{cases}$ (35)
@@
$\frac{d^{2}\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{y}_{s}d\mathbf{y}_{k}}=\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{s}}+\delta_{ks}\sum_{j=1}^{N}\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}\Pi_{kj}^{*}.$ (36)
@@
$\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{s}}=\sum_{j=1}^{N}\frac{\Pi^{*}_{kj}}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\left(\frac{\partial f^{*}_{k}}{\partial\mathbf{y}_{s}}+\frac{\partial g^{*}_{j}}{\partial\mathbf{y}_{s}}\right)$
$=\left[\begin{array}[]{ccccccccc}\bm{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]^{\top}\frac{(\mathbf{H}^{*})^{\dagger}}{\epsilon}\left[\begin{array}[]{ccccccccc}\bm{e}_{s}(\mathbf{B}_{s}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{s}^{\top}\end{array}\right],$
where we use the fact that $\left[\begin{array}[]{ccccccccc}\frac{\partial\mathbf{f}^{*}}{\partial\mathbf{y}_{s}}\\ \frac{\partial\mathbf{g}^{*}}{\partial\mathbf{y}_{s}}\end{array}\right]$ is the solution of the linear system (24), so it can be expressed as $\left[\begin{array}[]{ccccccccc}\frac{\partial\mathbf{f}^{*}}{\partial\mathbf{y}_{s}}\\ \frac{\partial\mathbf{g}^{*}}{\partial\mathbf{y}_{s}}\end{array}\right]=(\mathbf{H}^{*})^{\dagger}\left[\begin{array}[]{ccccccccc}\mathbf{e}_{s}(\mathbf{B}_{s}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{s}^{\top}\end{array}\right]$. For $k=s$, the term is expanded as follows
$\sum_{j=1}^{N}\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}\Pi_{kj}^{*}+\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{k}}$
$\,=\sum_{j=1}^{N}\frac{\Pi^{*}_{kj}}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\left(\frac{\partial f^{*}_{k}}{\partial\mathbf{y}_{k}}+\frac{\partial g^{*}_{j}}{\partial\mathbf{y}_{k}}\right)+\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}-\frac{1}{\epsilon}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)\right)$
$\,=\left[\begin{array}[]{ccccccccc}\bm{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]^{\top}\frac{(\mathbf{H}^{*})^{\dagger}}{\epsilon}\left[\begin{array}[]{ccccccccc}\bm{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]+\sum_{j=1}^{N}\Pi_{kj}^{*}\left(\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}-\frac{(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}})^{\top}(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}})}{\epsilon}\right).$
@@
Given the cost matrix $\mathbf{C}$ and the optimal coupling matrix $\bm{\Pi}^{*}$ from Sinkhorn algorithm, the Hession tensor is analytically calculated once the large linear system (32) is numerically solved. Unlike the previous approaches, the solution of the linear system is no longer needed to be differentiated in order to obtain the second-order derivatives, so our direct analytical Hessian expression could significantly speed up the computation with less memory burden. Similar to the case of the first-order derivative, the analytical expression for the Hessian is generic and applicable to any choice of cost function. It does not require $\bm{\mu}$ and $\bm{\nu}$ to be uniform either. If the Sinkhorn iterations stops early with $\widehat{\bm{\Pi}}$ being suboptimal, then the Hessian expression (34) is still exact for $\frac{d^{2}\mathrm{OT}_{\epsilon}(\mathbf{C},{\tilde{\bm{\mu}}},{\tilde{\bm{\nu}}})}{d\mathbf{Y}^{2}}$ with suboptimal coupling matrix $\widehat{\bm{\Pi}}$ and associated marginals $\tilde{\bm{\mu}}$ and $\tilde{\bm{\nu}}$, where $\sum_{j}\widehat{\Pi}_{ij}=\widetilde{\mu}_{i}$ and $\sum_{i}\widehat{\Pi}_{ij}=\widetilde{\nu}_{j}$.
With the analytical expression of Hessian, the following result show that the sum of the first index of the Hessian tensor is only dependent on the marginal probability vector $\bm{\mu}$. Later, it will be used as a marginal error to verify the accuracy of Hessian in the numerical implementation.
@@
If $C_{kj}=\|\mathbf{y}_{k}-\mathbf{y}_{j}^{*}\|_{2}^{2}$ for each $k,j$, then $\sum_{k=1}^{M}\mathcal{T}_{k\cdot s\cdot}=2\mu_{s}\mathbb{I}_{d}$, where $\mathbb{I}_{d}$ is the identity matrix of size $d$.
@@
Since the cost is square distance $C_{kj}=\|\mathbf{y}_{k}-\mathbf{y}_{j}^{*}\|_{2}^{2}$, then $\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}=2(\mathbf{y}_{k}-\mathbf{y}_{j}^{*})$ and $\frac{\partial^{2}C_{kj}}{\partial\mathbf{y}_{k}^{2}}=2\mathbb{I}_{d}$. The first term in the second-order derivative (36) expands as
$\sum_{k=1}^{M}\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\right)^{\top}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{s}}$ $=2\sum_{k=1}^{M}\mathbf{y}_{k}^{\top}\sum_{j=1}^{N}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{s}}-2\sum_{j=1}^{N}(\mathbf{y}_{j}^{*})^{\top}\sum_{k=1}^{M}\frac{\partial\Pi_{kj}^{*}}{\partial\mathbf{y}_{s}}=\mathbb{0}_{d\times d}.$
The last equal sign is because that $\sum_{j}\Pi^{*}_{kj}=\mu_{k}$ and $\sum_{k}\Pi^{*}_{kj}=\nu_{j}$, as well as $\sum_{j}\left(\frac{\partial\Pi^{*}_{kj}}{\partial\mathbf{y}_{s}}\right)=\frac{\partial\sum_{j}\Pi^{*}_{kj}}{\partial\mathbf{y}_{s}}$ and $\sum_{k}\left(\frac{\partial\Pi^{*}_{kj}}{\partial\mathbf{y}_{s}}\right)=\frac{\partial\sum_{k}\Pi^{*}_{kj}}{\partial\mathbf{y}_{s}}$.
The remaining term is $\sum_{j=1}^{N}\frac{\partial^{2}C_{sj}}{\partial\mathbf{y}_{s}^{2}}\Pi^{*}_{sj}=2\mathbb{I}_{d}\sum_{j=1}^{N}\Pi^{*}_{sj}=2\mu_{s}\mathbb{I}_{d}.$
@@ 3.4 Solve the linear systems with truncated SVD
A major challenge in computing the gradient of Sinkhorn distance in (33) and the Hessians of Entropy-regularized distance in (34) is that pseudo-inverse $(\mathbf{H}^{*})^{\dagger}$ can severely amplify the rounding errors or early stopping errors when the matrix $\mathbf{H}^{*}$ is ill-conditioned. The pseudo-inverse comes from the solutions to the linear systems (24) and (32) for computation of first- and second-order derivatives. Thus, instead of directly using $(\mathbf{H}^{*})^{\dagger}$, it is important to study the spectrum of the matrix $\mathbf{H}^{*}$ and regularize properly when the linear systems are ill-posed.
@@
Analytical and empirical results in the next section show that the $\mathbf{H}$-matrix is often ill-conditioned when the Sinkhorn regularization parameter $\epsilon$ is small or when the optimal coupling matrix $\bm{\Pi}^{*}$ is close to a permutation. Such ill-conditioned $\mathbf{H}$-matrices result in numerically unstable solutions when $\bm{\Pi}^{*}$ is slightly perturbed to $\widehat{\bm{\Pi}}$ due to the early stopping of the Sinkhorn iterations. This is also the exact reason that the previous implicit differentiation approach failed due to numerical instability.
@@
We tackle the potential ill-posedness by truncated Singular value decomposition (TSVD). We truncate the H-matrix’s spectrum up to the $K$-th largest eigenvalue, with $K=\max\{j:\frac{\lambda_{j}}{\lambda_{1}}>\alpha\}$, where $\{\lambda_{j}\}_{j=1}^{M+N}$ are the eigenvalues of $\mathbf{H}^{*}$ in descending order. In practice, we use the LAPACK’s DGELSD algorithm building in the least-square solver [4] and set $\alpha=10^{-10}$. The algorithm for gradient and Hessian computation is summarized in Algorithm 1.
@@
Algorithm 1 Computation of gradient and Hessian of EOT distance $\text{OT}_{\epsilon}(C,\bm{\mu},\bm{\nu})$ with respect to the source data $\bm{Y}$.
@@
Input: Optimized $\Pi^{*}$, cost matrix $C$, entropic regularization strength $\epsilon$, source data $\bm{Y}$ and singular value threshold $\alpha$.
@@
Output: Gradient with respect to $\bm{Y}$: $\frac{d\text{OT}_{\epsilon}(C,\bm{\mu},\bm{\nu})}{d\bm{Y}}\in\mathbb{R}^{M\times d}$, Hessian with respect to $\bm{Y}$: $\mathcal{T}\in\mathbb{R}^{M\times d\times M\times d}$
1:Compute the marginal probability vector $\bm{\mu}\in\mathbb{R}^{M}$ and $\bm{\nu}\in\mathbb{R}^{N}$: $\bm{\mu}\leftarrow\Pi^{*}\mathbb{1}_{N}$ and $\bm{\nu}\leftarrow(\Pi^{*})^{\top}\mathbb{1}_{M}$.
@@
2:Compute matrix $H\in\mathbb{R}^{(M+N)\times(M+N)}$: $H\leftarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\bm{\mu})&\Pi^{*}\\ (\Pi^{*})^{\top}&\text{diag}(\bm{\nu})\end{array}\right]$.
@@
3:Compute third-order tensor $\mathcal{B}\in\mathbb{R}^{M\times d\times N}$: $\mathcal{B}_{ksj}\leftarrow\frac{\partial C_{kj}}{\partial(\bm{y}_{k})_{s}}\Pi^{*}_{kj}$, compute third order tensor $\mathcal{R}\in\mathbb{R}^{(M+N)\times M\times d}$: $\mathcal{R}\leftarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbb{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]$ and compute the fourth-order tensor $\mathcal{E}$ in (35).
@@
4:Compute the gradient: $\frac{d\text{OT}_{\epsilon}(C,\bm{\mu},\bm{\nu})}{d\bm{Y}}=\mathcal{B}\cdot\mathbb{1}_{N}$.
@@
6:Approximate $\left[\begin{array}[]{ccccccccc}\frac{d\bm{f}^{*}}{d\bm{Y}}\\ \frac{d\bm{g}^{*}}{d\bm{Y}}\end{array}\right]$: $\left[\begin{array}[]{ccccccccc}\frac{d\bm{f}^{*}}{d\bm{Y}}\\ \frac{d\bm{g}^{*}}{d\bm{Y}}\end{array}\right]\leftarrow U_{K}\Lambda_{K}^{-1}U_{K}^{\top}\mathcal{R}$.
7:Compute the Hessian $\mathcal{T}$: $\mathcal{T}_{ktsl}\leftarrow\frac{1}{\epsilon}\sum_{i=1}^{M+N}\mathcal{R}_{ikt}\left[\begin{array}[]{ccccccccc}\frac{d\bm{f}^{*}}{d\bm{Y}}\\ \frac{d\bm{g}^{*}}{d\bm{Y}}\end{array}\right]_{isl}+\mathcal{E}_{ktsl}$.